OpenStudy (anonymous):

The solid is similar to a larger solid with the given scale factor. Find the surface area S and volume V of the larger solid S=28{pie}cm2 V=20{pie}cm3 Scale Factor-1:2

OpenStudy (anonymous):

please help me!!

OpenStudy (zepp):

Scale factor is 1 for 2, so if the larger solid has a side a 5, the large solid would have 10. but the area would be different, let's say the scale factor is k, then the area scale factor would be \(k^2\) and the volume scale factor would be \(k^3\)

OpenStudy (zepp):

And it's 'pi'

OpenStudy (zepp):

Thus, Area scale factor is 2^2 = 4 Volume scale factor is 2^3 = 8 \(S=28 \pi cm^2 * 4 = 112 cm^2 V=20 \pi cm^3 * 8= 160 cm^3\)